OpenStudy (anonymous):

Solve the equation (Tricky Equation!) Separable! (x+xy^2)dx+e^x^2 ydy=0

OpenStudy (anonymous):

\[(x+xy ^{2})dx+e ^{x ^{2}}ydy=0\]

OpenStudy (anonymous):

x(1 + y^2) dx = - (e^x^2) y dy -x e^(-x^2) dx = y/(1+y^2) dy integrating -1/2 e^(-x^2) = 1/2 ln(1 + y^2) + C -e^(-x^2) = ln(1 + y^2) + C ln(1 + y^2) = C - e^(-x^2) because the constant is still arbitrary y^2 + 1 = e^[C - e^(-x^2)] and thus you can generate two possible functions of y in terms of x soon.

OpenStudy (anonymous):

okay THANKS! give me 100000 seconds to process that please

OpenStudy (anonymous):

You're welcome!! Xo!